### v1

# Convex Physics Informed Neural Networks for the Monge-Ampère Optimal Transport Problem

Abstract.
Optimal transportation of raw material from suppliers to customers is an issue arising in logistics that is addressed here with a continuous model relying on optimal transport theory. A physics informed neural network method is advocated here for the solution of the corresponding generalized Monge-Ampère equation. Convex neural networks are advocated to enforce the convexity of the solution to the Monge-Ampère equation and obtain a suitable approximation of the optimal transport map. A particular focus is set on the enforcement of transport boundary conditions in the loss function. Numerical experiments illustrate the solution to the optimal transport problem in several configurations, and sensitivity analyses are performed.
Key words and phrases:
Optimal transport, Monge-Ampère equation, Convex PINNs, transport boundary conditions, Sensitivity analysis.
1. Introduction
Submitted to Engineering Computations on September 26, 2024
In the spirit of Monge [20], the optimal transport problem describes the problem of moving bulk material from a source (supplier) to a target destination location (customer), see, e.g., [7, 22, 23].
From the numerical viewpoint, many numerical methods from classical scientific computing have been advocated in the literature, involving, e.g., finite differences approaches [3, 5, 10] and finite element methods [4, 25, 30]. On the other hand, the emergence of deep neural networks has opened many avenues of investigation to approximate partial differential equations, see, e.g., [13, 16, 18]. In particular, Physics Informed Neural Networks (PINNs) have shown their capabilities to approximate solutions of partial differential equations [26]. Several extensions of PINNs are currently being investigated, see, e.g., consistent PINNs [6] or variational PINNS [17]. Applications of PINNs to the classical Monge-Ampère equation (with Dirichlet boundary conditions) can be found in [21], while a first tentative to solve the optimal transport problem is presented in [28].
We aim here at solving the following optimal transport problem. Let $\mathcal{X}$ and $\mathcal{Y}$ be two bounded, smooth open sets in $\mathbb{R}^{2}$, equipped respectively with probability measures $\mu(dx)=f(x)\,dx$ and $\nu(dy)=g(y)\,dy$, where the density functions $f$ and $g$ are supported on $\mathcal{X}$ and $\mathcal{Y}$, respectively. We further assume that $f$ and $g$ are strictly positive and bounded on their respective supports. Brenier’s Theorem, see, e.g., [14], implies that the optimal transport map, with respect to a quadratic cost, is given by the gradient of a convex potential $u$, which satisfies the generalized Monge-Ampère equation: find $u:\mathcal{X}\rightarrow\mathbb{R}$ such that
$\text{det}(D^{2}u(x))=\frac{f(x)}{g(\nabla u(x))}\quad\text{for }\mu\text{-almost every }x\in\mathcal{X},$
with the boundary condition
$\nabla u(\mathcal{X})=\mathcal{Y}.$
To approximate the optimal transport map, it is important to ensure that the numerical method provides an approximation of $u$ that is also convex. Otherwise, there is no guarantee that the solution is unique. To guarantee that a neural network approximation of the solution remains convex, so-called Input Convex Neural Networks have been designed in the literature [2]. First results for ICNNs applied to the optimal transport problem have been reported in [28]; they are specifically applied to the context of random sampling and density estimation based on available data, but do not aim at solving explicitly the partial differential equation. To ensure that the neural network approximation is convex, they require the activation function to be convex and non-decreasing, and all coefficients of the network to remain positive.
In this work, we propose to use ICNNs to approximate the solution of the generalized Monge-Ampère equation with transport boundary conditions. To enforce these boundary conditions at collocation points on the boundary, we incorporate a discretized version of the transport boundary condition into the loss function, inspired by the Hausdorff distance.
The structure of this article is as follows: Section 2 describes the derivation of the generalized Monge-Ampère equation. Section 3 describes the approach based on physics informed neural networks (PINNs), with, in particular ICNNs. Numerical experiments are reported in Section 4 to validate the approach and highlight the role of input convex PINNs for this problem.
2. Mathematical Model
The optimal transport problem, initially introduced by Monge [20], is formulated as follows: given two probability measures $\mu$ and $\nu$ defined on the measurable spaces $\mathcal{X}$ and $\mathcal{Y}$, find a measurable map $\mathbf{m}:\mathcal{X}\rightarrow\mathcal{Y}$ such that $\mathbf{m}_{\#}\mu=\nu$, i.e.,
$\nu(A)=\mu\left(\mathbf{m}^{-1}(A)\right)\quad\forall\,A\subset\mathcal{Y}\text{ measurable},$ (2.1)
in such a way that $\mathbf{m}$ minimizes the transportation cost:
[…]
where $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is the quadratic cost function $c(x,y)=\frac{1}{2}|x-y|^{2}$, and the minimum is taken over all measurable maps $\mathbf{s}:\mathcal{X}\rightarrow\mathcal{Y}$ such that $\mathbf{s}_{\#}\mu=\nu$. When the condition $\mathbf{m}_{\#}\mu=\nu$ is met, $\mathbf{m}$ is called a transport map, and if $\mathbf{m}$ also minimizes the cost, it is referred to as an optimal transport map. Figure 1 provides a visual sketch of this scenario.
$\Omega$$x$$\mathcal{X}$$\partial\mathcal{X}$$\mathcal{Y}$$\partial\mathcal{Y}$$\mathbf{m}(x)$
Figure 1. Sketch of the optimal transport problem, illustrating the mapping of a domain $\mathcal{X}$ onto another domain $\mathcal{Y}$.
Even for the quadratic cost, determining the existence of an optimal transport map is not straightforward. Additionally, examples can be constructed where the Monge problem is ill-posed due to the non-existence of a transport map; for instance, when $\mu$ is a Dirac measure and $\nu$ is not. This suggests that certain conditions on the measures $\mu$ and $\nu$ are necessary. Consider the case where $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{n}$, with $\mu(dx)=f(x)\,dx$ and $\nu(dy)=g(y)\,dy$. If $\mathbf{s}:\mathcal{X}\rightarrow\mathcal{Y}$ is a sufficiently smooth transport map, the condition $\mathbf{s}_{\#}\mu=\nu$ can be reformulated as a Jacobian equation. Specifically, if $\chi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a test function, the condition in (2.1) leads to
[…]
Assuming $\mathbf{s}$ is a diffeomorphism, we can set $y=\mathbf{s}(x)$ and apply a change of variables that yields
[…]
By the arbitrariness of $\chi$, this yields the Jacobian equation: $\mathbf{s}:\mathcal{X}\rightarrow\mathbb{R}^{n}$ such that
$f(x)=g(\mathbf{s}(x))\det(\nabla\mathbf{s}(x))\quad\text{a.e.}\ x\in\mathcal{X}.$
In the specific case where $\mu$ and $\nu$ are absolutely continuous probability measures with compact support in $\mathcal{X}$ and $\mathcal{Y}$, respectively, with density functions $f$ and $g$ that are bounded away from zero and infinity on $\mathcal{X}$ and $\mathcal{Y}$, respectively, Brenier’s theorem states that:
  1. i)
    There exists a unique solution $\mathbf{m}$ to the optimal transport problem (2.2).
  2. ii)
    There exists a convex function $u:\mathcal{X}\rightarrow\mathbb{R}$ such that the optimal map $\mathbf{m}$ is given by $\mathbf{m}(x)=\nabla u(x)$ for $\mu$-almost every $x$, and satisfies
    $\text{det}(D^{2}u(x))=\frac{f(x)}{g(\nabla u(x))}\quad\text{for }\mu\text{-a.e. }x\in\mathbb{R}^{n}.$
Moreover, the condition in (2.1) implies that the boundary condition $\nabla u(\mathcal{X})=\mathcal{Y}$ can be replaced by the boundary condition
$\nabla u(\partial\mathcal{X})=\partial\mathcal{Y}.$ (2.3)
In the sequel, let us now address the approximation of the optimal transport problem with neural networks.
3. Physics Informed Neural Networks
3.1. Generalities
We consider feedforward neural networks made up of an input layer, an output layer and $L\geq 1$ hidden layers. We denote by $N_{j}$ the number of neurons of the $j^{th}$ layer, $j=0,\ldots,L+1$. The network is parametrised by the weight matrices $W^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ and bias vectors $b^{(l)}\in\mathbb{R}^{N_{l+1}}$ with $l=0,\ldots,L$. The output is defined with respect to the input as:
$x^{0}=x,$
$x^{l}=\sigma(W^{(l-1)}x^{l-1}+b^{(l-1)}),\quad 1\leq l\leq L$
$y=W^{(L)}x^{L}+b^{(L)},$
where $\sigma$ is the activation function. The set of parameters is collectively denoted by $\theta=\big{\{}W^{(l)},b^{(l)}\big{\}}_{l}$, and the neural network mapping is denoted by $y:=u_{NN}(x)$. Common choices for $\sigma$ include the ReLU function, $\sigma(z)=\max(0,z)$, the hyperbolic tangent function, $\sigma(z)=\frac{e^{z}-e^{-z}}{e^{z}+e^{-z}}$, or the softplus function $\sigma(z)=\log(1+e^{z})$.
A neural network has the potential to provide an approximation $u_{NN}:\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L+1}}$ to a given mapping $u:\mathbb{R}^{N_{0}}\to\mathbb{R}^{N_{L+1}}$. Indeed, the universal approximation theorem [15] shows that, for single-layered neural networks and regardless of the shape of $\sigma$, this approximation is possible by density of the approximation spaces when the number $n$ of hidden neurons increases. Universal approximation theorems also apply to deep neural networks [15, 16] under some smoothness conditions.
3.2. Input Convex Neural Networks
As discussed in Section 2, the solution $u$ to the generalized Monge-Ampère equation must be a convex function. Therefore, it is imperative to ensure the convexity of the approximate solution $u_{NN}$. To this end, one can employ Input Convex Neural Networks (ICNNs), as introduced in [2], whose architecture inherently guarantees convexity with respect to the input. The general structure of an ICNN with $L$ hidden layers and input variables $x$ is given by
$x^{0}$ $=x,$ (3.1)
$x^{1}$ $=\sigma(L^{(0)}x^{0}+b^{(0)}),$
$x^{l}$ $=\sigma(W^{(l-1)}x^{l-1}+L^{(l-1)}x^{0}+b^{(l-1)}),\quad 2\leq l\leq L,$
$y$ $=W^{(L)}x^{L}+L^{(L)}x^{0}+b^{(L)}.$
Here, $\sigma$ denotes a specified increasing convex function, $W^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{l}}$ are matrices with nonnegative entries, $L^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{0}}$ are matrices, and $b^{(l)}\in\mathbb{R}^{N_{l+1}}$ are vectors. The convexity of the network follows from the following principles:
  1. i)
    A linear combination of convex functions with positive coefficients remains convex.
  2. ii)
    The composition $h\circ g$ of a convex function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}$ with an increasing convex function $h:\mathbb{R}\rightarrow\mathbb{R}$ is also convex.
A notable feature of ICNNs is the inclusion of passthrough layers, which directly connect the input $x$ to the hidden units in deeper layers. In traditional feedforward networks, such layers are unnecessary, as previous hidden units can be mapped to subsequent hidden units via an identity mapping. However, in ICNNs, the non-negativity constraint on the weights $W^{(i)}$ restricts the use of hidden units that replicate the identity mapping. Thus, this additional passthrough is explicitly incorporated to overcome this limitation and is crucial for the representational capacity of the network [11].
3.3. Physics Informed Neural Networks
To solve Monge-Ampère equations using neural networks approximation, we use Physics-Informed Neural Networks (PINNs) based on the strong formulation of the partial differential equation [26]. Consider the Monge-Ampère problem with Dirichlet boundary conditions: find $u:\mathcal{X}\rightarrow\mathbb{R}$ such that:
$\left\{\begin{array}[]{ll}\det{{D}^{2}u}=\dfrac{f(x)}{g(\nabla u(x))}&x\in\mathcal{X},\\ u(x)=h(x)&x\in\partial\mathcal{X},\end{array}\right.$ (3.2)
where $h$ is a prescribed function. The PINNs framework seeks a neural network approximation $u_{NN}$ of the solution $u$ to (3.2) by minimizing a weighted sum of the residual $\mathcal{L}_{PDE}$ and the boundary loss $\mathcal{L}_{b}$, specifically:
$u_{NN}:=\operatorname*{arg\,min}_{\theta}\mathcal{L}(\theta)=\operatorname*{arg\,min}_{\theta}\left\{\mathcal{L}_{PDE}(\theta)+C\mathcal{L}_{b}(\theta)\right\},$
where
$\mathcal{L}_{PDE}$ $=\left\|\det{D^{2}u_{NN}}-\displaystyle\frac{f}{g(\nabla u_{NN})}\right\|_{L^{2}(\mathcal{X})},$
$\mathcal{L}_{b}$ $=\|u_{NN}-h\|_{L^{2}(\partial\mathcal{X})},$
and $C>0$ controls the weight of the boundary conditions. To numerically approximate $\mathcal{L}_{PDE}$ and $\mathcal{L}_{b}$, we select a set of collocation points $\{{p}_{i}\}_{i=1}^{N_{c}}\subset\mathcal{X}$ and boundary points $\{{q}_{i}\}_{i=1}^{N_{b}}\subset\partial\mathcal{X}$, and approximate the $L^{2}$ norms as follows:
$\mathcal{L}_{PDE}\approx$ $E_{PDE}:=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}\left|\det{{D}^{2}u_{NN}({p}_{i})}-\frac{f({p}_{i})}{g(\nabla u_{NN}({p}_{i}))}\right|^{2},$
$\mathcal{L}_{b}\approx$ $E_{b}:=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}|u_{NN}({q}_{i})-h({q}_{i})|^{2}.$
When the Dirichlet boundary conditions in (3.2) are replaced by the transport boundary conditions (2.3), we consider the problem
$\left\{\begin{array}[]{ll}\det{{D}^{2}u}=\dfrac{f(x)}{g(\nabla u(x))}&x\in\mathcal{X},\\ \nabla u(\partial\mathcal{X})=\partial\mathcal{Y}.\end{array}\right.$ (3.3)
The residual $E_{b}$ must be adjusted accordingly. To do so, we notice that
$\nabla u(\partial\mathcal{X})=\partial\mathcal{Y}\iff d_{H}(\nabla u(\partial\mathcal{X}),\partial\mathcal{Y})=0,$
where $d_{H}$ is the Hausdorff distance with respect to the Euclidean metric, which is defined by
$d_{H}(A,B):=\max\left\{\sup_{a\in A}\mathrm{dist}(a,B),\;\sup_{b\in B}\mathrm{dist}(A,b)\right\},\quad A,B\subset\mathbb{R}^{n},$
with $\mathrm{dist}(a,B):=\min_{b\in B}\|a-b\|_{2}$. The Hausdorff distance not only ensures that a point in $\partial\mathcal{X}$ is mapped through $\nabla u$ to a point in $\partial\mathcal{Y}$ but also guarantees the surjectivity of the map $\nabla u$. On a discrete level we approximate $d_{H}$ in the following way: suppose we have two sets of boundary points
$\{{x_{i}}\}_{i=1}^{N_{bx}}\subset\partial\mathcal{X},\quad\{{y_{i}}\}_{i=1}^{N_{by}}\subset\partial\mathcal{Y},$
the boundary loss $E_{OT}$ is then defined as
$E_{OT}=\frac{1}{N_{bx}}\sum_{i=1}^{N_{bx}}\mathrm{dist}\left(\nabla u({x_{i}}),\{{y_{j}}\}_{j=1}^{N_{by}}\right)^{2}+\frac{1}{N_{by}}\sum_{i=1}^{N_{by}}\mathrm{dist}\left(\nabla u\big{(}\{{x_{j}}\}_{j=1}^{N_{bx}}\big{)},{y_{i}}\right)^{2},$ (3.4)
where the first term in (3.4) controls the injectivity of the map $\nabla u_{NN}$ while the second term its surjectivity. Finally, we look for $u_{NN}$ convex such that
$u_{NN}=\operatorname*{arg\,min}_{\theta}\left\{E_{PDE}+CE_{OT}\right\}.$
4. Numerical experiments
Several numerical experiments are presented in order to validate the convergence and accuracy properties of the convex PINNs methodology.
4.1. Implementation details
For the implementation of the PINN solver, the code has been developed in pytorch and the training is performed using input convex neural networks (ICNNs). The neural networks used in the test cases consist of 4 hidden layers with 10 neurons per layer. The activation function employed is the softplus function, defined as $\sigma(x)=\log(1+e^{x})$, chosen because it is an increasing convex function. To guarantee the positivity of the weight matrices $W^{(l)}\in\mathbb{R}^{N_{l+1}\times N_{l}}$, the weights are initialized by squaring them element-wise, i.e., $w_{ij}\leftarrow w_{ij}^{2}$, as suggested in [21]. Note that the alternative approach of clipping the weights, such that $w_{ij}\geq 0$ for all $i,j$, following [2], does not provide satisfactory results in our experiments.
For the optimization method, we use the L-BFGS algorithm. However, given the sensitivity of L-BFGS to initialization, we first perform several iterations of the Adam optimizer to provide a more robust starting point. This approach allows for more stable convergence of the L-BFGS optimizer.
Moreover, ICNNs are notoriously difficult to train [8, 27] and a meaningful initialization is required. In particular, the network parameters are initialized by training the network to approximate the identity map for the gradient of the solution, $\nabla u\approx\text{Id}$, inspired by the work in [28]. The neural networs are trained for $2500$ L-BFGS epochs.
By default, every test case is run with $N_{c}=800$ collocation points on the interior of the domain and $N_{b}=800$ points on the boundary, choice which will be justified in the numerical experiments in the sequel. The constant $C$ is set to 1, unless specified otherwise.
4.2. Transport of a constant density from a unit disk domain to an ellipse domain
The first example aims at transporting a constant probability density function supported in a unit disk domain $\mathcal{X}\subset\mathbb{R}^{2}$ to another constant probability density function whose support is an ellipse domain $\mathcal{Y}\subset\mathbb{R}^{2}$. Let us thus define
$\mathcal{X}=\{x=(x_{1},x_{2})\in\mathbb{R}^{2}:\,x_{1}^{2}+x_{2}^{2}<1\},$
and
$\mathcal{Y}=\left\{y=(y_{1},y_{2})\in\mathbb{R}^{2}:\,\frac{\left(y_{1}-3.5\right)^{2}}{(2)^{2}}+\frac{y_{2}^{2}}{(0.5)^{2}}<1\right\}.$
The corresponding generalized Monge-Ampère problem (3.3) is considered here with $f(x)=\frac{1}{\pi}\chi_{\mathcal{X}}(x)$, $g(y)=\frac{1}{\pi}\chi_{\mathcal{Y}}(y)$, where $\chi_{A}$ denotes the characteristic function of the domain $A$. With this data, the exact solution to (3.3) is
$\nabla u_{ex}(x)=\begin{pmatrix}2x_{1}+\frac{7}{2}\\ \frac{1}{2}x_{2}\end{pmatrix},\quad x=(x_{1},x_{2})\in\mathcal{X}.$
Figure 2 shows how uniformly sampled points in the unit ball are transported to the ellipse, both with the exact solution and the approximate map. It illustrates an appropriate, uniform, repartition of the sampled points (left), and an accurate transport of those points, whose final position coincides well with their exact values (right). Figure 3 visualizes the corresponding vector field $\nabla u_{NN}$, while Figure 4 details the first and second components of $\nabla u_{NN}$, together with the numerical error component-wise. One can observe that the error on both components is well balanced and of the order $10^{-4}$.
Figure 2. Transport of a constant density from a unit disk domain to an ellipse domain. Visualization of uniformly sampled source points and their image through exact and approximate transport maps.
Figure 3. Transport of a constant density from a unit disk domain to an ellipse domain. Visualization of the transport map $\nabla u_{NN}$.
Figure 4. Transport of a constant density from a unit disk domain to an ellipse domain. Top left: first $x$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,x}-\nabla u_{ex,x}\right|$; Bottom left: second $y$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,y}-\nabla u_{ex,y}\right|$.
Let us study the sensitivity of those results with respect to the number of epochs and number of collocation points. Figure 5 (top and middle) visualizes the behavior of the $L^{2}$ error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ with respect to these factors respectively. The error has been averaged over a sample of $10$ simulations. In particular, when the number of collocation points increases, e.g., from $N_{c}=50$ to $N_{c}=100$, the first $50$ points in each sample remains the same as those for the smaller sample. It shows that the error decreases when either of these quantities increases, albeit with a large sample variation. Finally, Figure 5 (bottom) visualizes the behavior of the $L^{2}$ error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ with respect to the ratio $N_{b}/N_{c}$ between the number of boundary points and the number of collocation points. It shows that the error is smaller for a ratio equal to one, which confirms the initial choice of $N_{b}=N_{c}$ made earlier.
Figure 5. Transport of a constant density from a unit disk domain to an ellipse domain. Top: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs number of epochs. Middle: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs number of collocation points. Bottom: $L^{2}$ (test) error $\left|\left|\nabla u_{ex}-\nabla u_{NN}\right|\right|_{L^{2}(\mathcal{X})}$ vs ratio $N_{b}/N_{c}$. (Results averaged on $10$ simulations.)
4.3. Transport of a constant density from an ellipse domain to a rotated ellipse domain
The second example aims at transporting a constant probability density function supported in an ellipse domain $\mathcal{X}\subset\mathbb{R}^{2}$ to another constant probability density function whose support is a rotated ellipse domain $\mathcal{Y}\subset\mathbb{R}^{2}$. More precisely, let $\mathcal{B}_{1}$ be the unit disk in $\mathbb{R}^{2}$ and $\mathcal{X}=M_{\mathcal{X}}\mathcal{B}_{1}$ and $\mathcal{Y}=M_{\mathcal{Y}}\mathcal{B}_{1}$, $M_{\mathcal{X}},M_{\mathcal{Y}}$ being positive definite matrices. The optimal map can be explicitly calculated in $\mathbb{R}^{2}$ and is given by
$T({x})=M_{\mathcal{Y}}\mathcal{R}_{\theta}M_{\mathcal{X}}^{-1}{x},\qquad{x}\in\mathcal{X},$
where $\mathcal{R}_{\theta}$ is the rotation matrix
$\mathcal{R}_{\theta}=\begin{pmatrix}\cos(\theta)&-\sin(\theta)\\ \sin(\theta)&\cos(\theta)\end{pmatrix},$
and the angle $\theta$ is calculated as follows:
$\tan(\theta)=\mathrm{trace}(M_{\mathcal{X}}^{-1}M_{\mathcal{Y}}^{-1}J)/\mathrm{trace}(M_{\mathcal{X}}^{-1}M_{\mathcal{Y}}^{-1}),$
where
$J=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}.$
We test the following example where
$M_{\mathcal{X}}=\begin{pmatrix}0.8&0\\ 0&0.4\end{pmatrix},\qquad M_{\mathcal{Y}}=\begin{pmatrix}0.8&0.2\\ 0.2&0.6\end{pmatrix}.$
We run the optimising process with $400$ Adam epochs and $1000$ L-BFGS epochs. Figure 6 shows how uniformly sampled points in the original ellipse $\mathcal{X}$ are transported to the rotated ellipse $\mathcal{Y}$; it illustrates an appropriate, uniform, repartition of the sampled points and an accurate transport of those points, whose final position coincides well with their exact values. Figure 7 visualizes the corresponding vector field $\nabla u_{NN}$. Figure 8 visualizes the approximation error on both components of $\nabla u_{NN}$, and highlights that the error is of the order of $10^{-3}$ pointwise.
Figure 6. Transport of a constant density from an ellipse domain to a rotated ellipse domain. Visualization of uniformly sampled source points and their image through exact and approximate transport maps.
Figure 7. Transport of a constant density from an ellipse domain to a rotated ellipse domain Visualization of the transport map $\nabla u_{NN}$.
Figure 8. Transport of a constant density from a unit disk domain to an ellipse domain. Top left: first $x$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,x}-\nabla u_{ex,x}\right|$; Bottom left: second $y$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,y}-\nabla u_{ex,y}\right|$.
4.4. Transport of a Gaussian density in the unit square to a uniform density in the unit square
In order to mimic the optimal transport of piles of debris [20], we transport the unit square domain with an initial Gaussian density function, namely
$f({x})=c_{0}\exp\left\{-\dfrac{1}{2\sigma^{2}}({x}-{x}_{0})^{2}\right\}\chi_{[0,1]^{2}}({x}),$ (4.1)
where $c_{0}$ is a normalization constant, $\sigma^{2}=0.25$ and ${x}_{0}=(0.25,0.75)$. This function is transported into a uniform density on the same unit square domain:
$g({x})=\chi_{[0,1]^{2}}({x}).$
Figure 9 illustrates the approximated optimal transport map. The histogram in Figure 9 (left) shows the distribution of points sampled from the Gaussian distribution (4.1) in the unit square. Then the points are transported to a uniform distribution through the exact map $\nabla u$ (center) and the approximated map $\nabla u_{NN}$ (right).
We can observe that the transported solution re-dispatches appropriately uniformly the points on the grid, and that the exact and the approximated solutions are very similar to each other. Figure 10 visualizes the approximation error on both components of $\nabla u_{NN}$. The error is still of a similar order, and we can observe that it concentrates in the corners of the unit square domain.
Figure 9. Transport of a Gaussian density to a uniform density. Left: sampling of points according to Gaussian density (4.1); Middle: transported solution with exact solution $\nabla u_{ex}$; Right: transported solution with approximated solution $\nabla u_{NN}$. ($N=100000$ points sampled into the cells of a structured grid of size $80\times 80$).
Figure 10. Transport of a Gaussian density to a uniform density. Top left: first $x$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,x}-\nabla u_{ex,x}\right|$; Bottom left: second $y$-coordinate of the transport map $\nabla u_{NN}$; Top right: Pointwise absolute error $\left|\nabla u_{NN,y}-\nabla u_{ex,y}\right|$.
4.5. Transport of a Gaussian density in the unit square to another Gaussian density in the unit square
We consider the transport of an initial Gaussian density function within the unit square domain into another Gaussian density function, in order to evaluate the numerical diffusion of the method when reconstructing a non-constant target density function. More precisely, we consider:
$f({x})=c_{0}\exp\left\{-\dfrac{1}{2\sigma^{2}}({x}-{x}_{0})^{2}\right\}\chi_{[0,1]^{2}}({x}).$
where $c_{0}$ is a normalization constant, $\sigma^{2}=0.25$ and ${x}_{0}=(0.25,0.75)$, and $g({x})$ defined similarly, with $\sigma^{2}=0.25$ and ${x}_{0}=(0.75,0.25)$. The histogram in Figure 11 (left) shows the distribution of points sampled from the source Gaussian distribution (4.1) in the unit square and then transported to the target Gaussian distribution through the map $\nabla u_{NN}$ (right). Even if there is no analytical solution, we observe that the distribution of the transported points very closely resembles a Gaussian distribution.
Figure 11. Transport of a Gaussian density to another Gaussian density. Left: sampling of points according to original Gaussian density; Right: transported solution with approximated solution $\nabla u_{NN}$. ($N=100000$ points sampled into the cells of a structured grid of size $80\times 80$).
4.6. Transport of a bimodal Gaussian density in the unit square to a uniform density in the unit square
Finally, in an effort to mimic several suppliers transporting material from several sources, we present an example of multiple Gaussian sources being scattered into a uniform density field. We consider an anisotropic bimodal Gaussian distribution on the unit square, and a target uniform distribution on the same unit square, namely we define, with $x=(x_{1},x_{2})$:
$f({x})$ $=$ $c_{0}\left(\exp\left\{-\dfrac{1}{2\sigma^{2}_{xx}}(x_{1}-0.5)^{2}-\dfrac{1}{2\sigma^{2}_{yy}}(x_{2}-0.2)^{2}\right\}\right.$ (4.2)
$\left.+\exp\left\{-\dfrac{1}{2\sigma^{2}_{xx}}(x_{1}-0.5)^{2}-\dfrac{1}{2\sigma^{2}_{yy}}(x_{2}-0.8)^{2}\right\}\right)\chi_{[0,1]^{2}}({x}),$
where $c_{0}$ is a normalization constant, $\sigma^{2}_{xx}=0.25$ and $\sigma^{2}_{yy}=0.015625$, and
$g({x})=\chi_{[0,1]^{2}}({x}).$
The histogram in Figure 12 (left) shows the distribution of points sampled from the source bimodal Gaussian distribution (4.2) in the unit square and then transported to the uniform distribution through the map $\nabla u_{NN}$ (right). Even if the exact solution $u$ is unknown, we observe that the distribution of the transported points is uniform.
Figure 12. Transport of a bimodal Gaussian density to a uniform density. Left: sampling of points according to bimodal Gaussian density (4.2); Right: transported solution with approximated solution $\nabla u_{NN}$. ($N=100000$ points sampled into the cells of a structured grid of size $80\times 80$).
4.7. Transport of a 3D Gaussian density in the unit cube to a uniform density in the unit square
Finally, to illustrate the ability of the algorithm to be effective in more than two space dimensions, the test case in Section 4.4 is extended to the 3D case. We consider
$f({x})=c_{0}\exp\left\{-\dfrac{1}{2\sigma^{2}}({x}-{x}_{0})^{2}\right\}\chi_{[0,1]^{3}}({x}),$ (4.3)
where $c_{0}$ is a normalization constant, $\sigma^{2}=0.125$ and ${x}_{0}=(0.75,0.75,0.75)$. This function is transported into a uniform density on the unit cube:
$g({x})=\chi_{[0,1]^{3}}({x}).$
Figure 13 illustrates the approximated optimal transport map. We can observe that the transported solution re-dispatches appropriately uniformly the points on the grid. Figure 14 illustrates the transport of uniformly randomly sampled boundary points. We observe that $\nabla u_{NN}(\partial\mathcal{X})=\partial\mathcal{Y}$ as well.
Figure 13. Transport of a 3D Gaussian density to a uniform density. Left: sampling of points inside the unit cube $\mathcal{X}$ according to Gaussian density (4.3); Right: location of the inside points transported with the approximated solution $\nabla u_{NN}$.
Figure 14. Transport of a 3D Gaussian density to a uniform density. Left: sample of points on the boundary of the unit cube $\partial\mathcal{X}$, uniformly sampled on the boundary Right: location of the boundary points transported with the approximated solution $\nabla u_{NN}$.
5. Conclusions and Perspectives
A model based on a generalized Monge-Ampère equation with transport boundary conditions has been considered to solve the optimal transport problem. A convex input PINNs method has been advocated for the numerical approximation of its convex solution. The loss function has been modified to account for the transport boundary conditions by using a discrete version of the Hausdorff distance.
Since we look for a convex solution of the Monge-Ampère equation, input convex neural networks are suitable to obtain appropriate approximations as they are seeking for solutions in the right subspaces. However, input convex neural networks are notoriously difficult to train.
A sensitivity analysis has shown that the model is not particularly affected by the random sampling of the collocation and boundary points. Numerical results have focused on transporting uniform and Gaussian distribution functions in two dimensions of space. However, we have shown that results can be extended to higher dimensions.
Perspectives include the further incorporation of several sources and targets, via multimodal Gaussian density functions, for real-life applications, as well as the treatment of obstacles. The investigation of the scaling up of the algorithm in a larger number of dimensions is also of interest.
Acknowledgements
The authors thank Archie Lawson (EPFL) for implementation support during an EPFL Bachelor semester project, and Prof. Marco Picasso (EPFL) for fruitful discussions.